# &Jiawei Li

### v1

@@ Appendix B Experiment Details
B.1 Predefined Set for Image Augmentations
@@
B.2 Attack Methods
The attack target and victim class are chosen at random for each trial. We shall now discuss the details for each attack method below. The relevant code is additionally publicly available in our repository: https://github.com/andyisokay/delta-influence.
BadNet For CIFAR datasets, we add a $3\times 3$ checkboard-patterned black patch (pixel values set to zero) at the bottom-right corner of each $32\times 32$ image. For the Imagenette dataset, we utilize a larger square $22\times 2$ 22 black patch to ensure successful injection of the poison. The number of poisoned images varies by dataset: 500 for CIFAR10, 350 for CIFAR100, and 858 for Imagenette.
@@ B.2 Attack Methods
B.3 Hyperparameters for Detection Methods
@@ Appendix C Results for Ablating Image-Only and Label-Only Augmentations
We show in Figure 7 that Label-Only augmentations are effective in removing the data poisoning (lower poison success rate), while Image-Only augmentations perform poorly in this regard. However, as demonstrated in Table 2, Label-Only augmentations lead to the unnecessary discard of many clean samples, whereas image augmentations significantly reduce the false positive rate, preserving clean data and improving detection precision. Therefore both label and image augmentations are crucial to the effectiveness of the $\Delta$-Influence method.

### v2

@@ Appendix B Experiment Details
B.1 Poisoned Training Sample Is Not a Reliable Target for Influence-Based Unlearning
Given that a small subset of poisoned training data—commonly referred to as a forget set (Goel et al., 2024), could be identified as the prerequisite for unlearning, a natural question arises: why $\Delta-\mathrm{Influence}$ focus on an identified affected test sample rather than simply using a poisoned training sample?
One overlooked fact is for some covert attack like Witches’ Brew, the attack pattern is different between training and testing, which is not the case for Frequency Trigger and BadNet. Moreover, the clean-label attack manner and the imperceptible perturbations make it notoriously difficult to identify training poisons for such attacks. However, we emphasize that, regardless of how clever and stealthy an attack is designed, its primary goal is to alter model predictions on specific test points, making anomalies more apparent after deployment. Hence we think that having one identified test point is generally more feasible than identifying a poisoned training point in this context and better suited for influence-based analysis, which attributes model behavior to particular training instances.
What’s more, as demonstrated in Table˜4, taking Witches’ Brew as an example, we find that even when defenders can reveal a poisoned training sample, the poisoned behavior cannot be reliably mitigated, while we show that $\Delta-\mathrm{Influence}$, utilizing an identified poisoned test point, can systematically undo the attack’s impact.
This underscores a fundamental limitation of how influence functions work: influence functions inherently rely on clear causal relationships, where specific training samples directly impact corresponding test-time anomalies. However, in poisoned learning scenarios, such causality is often obscured: while the training poisons as a whole shifts model behavior, it’s causal effect with one individual poisoned sample in it could be more ambiguous. This intuition that using a poisoned training sample as the target is less reliable than using an affected test point, is further supported by our empirical findings (as shown in Table˜4, influence-based methods fail to unlearn poisons when guided by a poisoned train point). We hope this observation provides useful insights for how target selection impacts causal tracing effectiveness in influence-based unlearning.
Based on the above observations, we suggest that using an identified poisoned test point for influence-based unlearning. Although when the attack pattern is consistent between training and testing (e.g. Frequency Trigger (Zeng et al., 2021) and BadNet (Gu et al., 2019)), using a poisoned training sample as the target also work, We argue that defenders should not assume such prior knowledge, e.g. what attack is performed and what the attack pattern is, which is rarely available in practice. Hence using a poisoned test point is more reliable and generalizable across different attack scenarios.
Finally, although it’s not the focus of this paper, here we discuss approaches to get such a test point, realistic scenarios include: (i) whitehat adversarial research teams conducting jailbreaking-style tests to expose failure modes; (ii) Companies internally systematically stress-testing for vulnerabilities; and (iii) Companies using anomaly detection algorithms to monitor user interactions for abnormal behavior. Note that determining whether a test point is harmful or benign relies on the developer’s domain expertise, this largely unexplored area is increasingly necessary due to massive training datasets and the rise of opaque open-source base models, offering promising directions for future research.
B.2 Predefined Set for Image Augmentations
@@
B.3 Attack Methods
The attack target and victim class are chosen at random for each trial. We shall now discuss the details for each attack method below. The relevant code is additionally publicly available in our repository: https://github.com/Ruby-a07/delta-influence.
BadNet For CIFAR datasets, we add a $3\times 3$ checkboard-patterned black patch (pixel values set to zero) at the bottom-right corner of each $32\times 32$ image. For the Imagenette dataset, we utilize a larger square $22\times 22$ black patch to ensure successful injection of the poison. The number of poisoned images varies by dataset: 500 for CIFAR10, 350 for CIFAR100, and 858 for Imagenette.
@@ B.3 Attack Methods
B.4 Hyperparameters for Detection Methods
@@ Appendix C Results for Ablating Image-Only and Label-Only Augmentations
We show in Figure˜6 that Label-Only augmentations are effective in removing the data poisoning (lower poison success rate), while Image-Only augmentations perform poorly in this regard. However, as demonstrated in Table˜2, Label-Only augmentations lead to the unnecessary discard of many clean samples, whereas image augmentations significantly reduce the false positive rate, preserving clean data and improving detection precision. Therefore both label and image augmentations are crucial to the effectiveness of the $\Delta$-Influence method.
Appendix D Does $\Delta$-Influence Perform the Best Across Unlearning Algorithms?
Results. The evaluation results in Figure˜7 shows that $\Delta-\mathrm{Influence}$ outperforms other methods, unlearning poisons in all six cases with minimal performance loss. In contrast, EK-FAC, ActClust, and SpecSig performed randomly, achieving unlearning primarily because even a randomly initialized model would not retain poisoning. Performance drops were primarily due to SCRUB’s sensitivity to false positives from its gradient ascent step. FreqDef avoided randomness but failed to unlearn poisons in all cases. Notably, $\Delta-\mathrm{Influence}$ minimized false positives, maintaining consistent and reliable outcomes.
Appendix E Limitations
$\Delta-\mathrm{Influence}$ is based on influence functions and hence inherit their drawbacks. Possible attacks like those are only injected during test phase can evade our detection.
Appendix F Compute Resources
We used a single NVIDIA A100-80GB GPU for all experiments. The training time for each experiment varied depending on the model, dataset, and hyperparameters. On ImageNette poisoned by Witches’ Brew (with the according hyperparameters provided in the appendix), the time of detection is normally half an hour. Note that this could be speed up by parallelizing the computation of influence scores but this is not the focus of this paper.